# Kairosis: A method for dynamical probability forecast aggregation informed by Bayesian change point detection

### v2

@@ Abstract
We present a new method, “kairosis”, for aggregating probability forecasts made over a time period of a single outcome determined at the end of that period. Informed by work on Bayesian change-point detection, we begin by constructing for each time during the period a posterior probability that the forecasts before and after this time are distributed differently. The distribution of these probabilities is then integrated to give a cumulative mass function, which is then used to calculate a weighted median forecast. The effect is to construct a pool in which those forecasts are most heavily weighted which have been made since the likely most recent change in the forecasts’ distribution. Kairosis outperforms standard methods, and is especially suitable for geopolitical forecasting tournaments because it is observed to be robust across disparate questions and forecaster distributions.
@@ 1 Introduction
Geopolitical forecasting tournaments have become increasingly popular over the last decade, notable providers including the Good Judgment Project and Metaculus. A typical question from Metaculus is that of Figure 1, “Will Donald Trump be president of the USA in 2019?”. From when the question opened (May 17, 2017), forecasters submitted probability forecasts (on a scale of 00 to $1$), until the question was resolved on Feb 1, 2019, although here we show only the first seven months’ forecasts. A forecaster may make as many or as few forecasts as they wish, at any time they wish; thus we observe only the overall evolving distribution of forecasts, rather than multiple time-series of individual forecasters’ evolving beliefs.
After resolution, the forecasts are scored. If forecasts are considered “static”, taking no account of when the forecast is submitted, a simple proper probability score, such as the Brier (quadratic, Brier 1950) or Log (logarithmic, Good 1952) score, can be used. Proper scores are optimized by, and therefore incentivize forecasters to submit their best estimates of, the true probability – although propriety fails if rewards are not proportional to the score, for example if the prize goes to the overall winner (see for example Roulston 2007, Broecker and Smith 2007). But prescience is clearly valuable, and Metaculus, for example, weights the score by how long it was submitted before resolution. Indeed it is clear just from a visual inspection that the distribution of forecasts is not stationary; it changes both smoothly and sharply at certain points, as does the density of forecasts submitted. The reason is obvious: news occurs much in this way, and new information is continually informing the forecasting process.
@@
However, as Himmelstein et al. (2023) note, “it is possible to envision a hybrid method … of the selection and weighting approaches”, and it is our purpose in this article to provide this. Our technique is informed by, but not equivalent to, Bayesian change point detection, and can be viewed as something like an interpolation between exponential discounting and the “most recent change point” method suggested above. For any given present time when the question is open it begins in a Bayesian manner with an exponential prior on the time of the most recent previous change point. We then separate the unit probability interval into a number of bins and model the number of forecasts in each one using a compound Dirichlet-categorical distribution. For each time $t$ between the question-opening and the present, we compute the likelihood that $t$ splits the forecasts into two sets of forecasts with distinct distributions. This gives us, for all past $t$, a set of posterior probabilities for the possible change point locations. The posterior mass is integrated to give a cumulative mass function (CMF), ranging from 0 at the opening time to 1 at the present, which is then used to weight past forecasts. Our final, aggregate forecast is then the CMF-weighted median (i.e. the median of the weighted distribution).
The practical effect of our method is to create a weighting with multiple downward (as we move into the past) steps, each corresponding to significant changes in the distribution of forecasts. At one extreme, if there is no obvious time at which the distribution of forecasts changes, the weighting stays close to the original exponential discounting of past forecasts. At the other extreme a single, obvious change point between very different distributions of forecasts is effectively a horizon, behind which old forecasts add nothing.
@@
It is clear from Figure 1 that forecasts are not made at a uniform rate in chronological time. We account for this by introducing “forecaster time” and a “forecaster clock” which increments forwards by one unit (i.e. it “ticks”) every time a forecast is received. Thus the transformation from calendar time to forecaster time effectively dilates and magnifies the moments at which forecasts are being made at a high rate, typically due to the publication of relevant news, and compresses periods of forecaster inactivity. Intense forecasting activity, however, does not necessarily imply that the information landscape and the distribution of beliefs among forecasters is changing. To describe the speed of this type of change we find it useful to introduce the notion of “kairos time” and the “kairos clock”. These terms are inspired by the ancient Greek $\kappa\alpha\iota\rho\grave{o}\zeta$ (kairos), referring to the “time” of critical moments of lived experience, which is distinguished from $\chi\rho\acute{o}\nu o\zeta$ (chronos), corresponding to our notion of chronological or “calendar” time (Smith, 1969). It is this kairos time that we ultimately want to use to discount older forecasts when aggregating them. In the relevant literature “kairos” has a nice interpretation as “a moment of time when a prophecy was pronounced” (Tzamalikos, 2007), which fits comfortably with our conjecture that change points in forecast distributions are being driven by the arrival of new information. “Kairosis” makes for a correspondingly concise name for our method, and in particular for the direct, nonlinear transformation from chronos to kairos effected by the CMF for putative change points (an example of which is presented in Figure 3).
@@
The article is structured as follows. In Section 2 we describe our method in detail, illustrating in Figures 13 its application to the question of the identity of the next US President. In Section 3, after establishing a set of metrics and comparators, we report empirical results across 650 forecasting problems posed my Metaculus. The main part of the paper finishes with a discussion of possible developments and applications. In A we provide an additional illustrative example of computing a forecast-weighting function with our kairosis methodology in order to cement the ideas presented in the body of the paper. In B we present the preliminary results of the extension of kairosis to point forecasts of continuous variables. In C we report the results of a sensitivity analysis that accompanies the findings of Section 3.
@@ 2.1 Deriving a distribution over change point locations
Now, returning to our central example, suppose it is October 10, 2017, and we are tasked with submitting an optimal aggregated probability forecast to answer the question of the 2019 US President. It is reasonable to assume that most of the data is relevant, but to what extent is not immediately apparent. One might, for example, propose that more recent data is likely to be more informative than that from the distant past. On the assumption that critical events invalidating preceding forecasts occur independently at a constant rate over time, we are led naturally to the idea of exponentially discounting old forecasts.
We therefore begin with a prior assumption of a constant probability $p\in[0,1]$ that at least one change point-inducing event occurs between consecutive forecasts. This motivates a geometric prior distribution on the time of the last change point, $P(t_{\rm\small CP}=t)=p(1-p)^{N-t}$ where $N$ is the total number of forecasts (up to the present time, here October 10, 2017), which follows from the idea that for $t$ to have been the last change point there must have been $N-t$ subsequent inter-forecast periods without a change point. We then update our prior distribution given knowledge of observed forecasts using Bayes’ theorem,
$P(t_{\rm\small CP}=t|\text{Forecasts})=\frac{{P(\text{Forecasts}|t_{\rm\small CP}=t)}{P(t_{\rm\small CP}=t)}}{P(\text{Forecasts})}.$ (3)
The next key ingredient for our methodology is thus to specify the distribution of the forecasts given a candidate change point. Inspired by work on Bayesian hypothesis testing by Holmes et al. (2015), we use the compound Dirichlet-categorical distribution to describe the number of forecasts falling in different sub-intervals (“bins”) of $[0,1]$. We use the probabilities this distribution assigns to the observed bin-counts to inform the quantity $P(\text{Forecasts}|t_{\rm\small CP}=t)$ appearing in (3).
@@
Now, to propose a change point at time $t_{\rm\small CP}=t$ is to suppose that forecasts before and after $t$ follow two independent Dirichlet-categorical distributions, because an event is thought to have occurred at $t$ that has fundamentally changed the forecasters’ (unobserved) distribution of beliefs. Of course this is never wholly true – forecasts always incorporate the previously-available information which informed forecasts before $t$ – but we now have a way of quantifying the probability that it is so. Thus, for a single change point at $t$, expression (3) becomes
$P(t_{\rm\small CP}=t|\text{Forecasts})\propto$ ${P(\text{Forecasts}|t_{\rm\small CP}=t)}{P(t_{\rm\small CP}=t)}$
@@
where the unprimed and primed letters denote counts $n$ and pseudo-counts $\alpha$ before and after $t$.
@@
We then evaluate (6) for every candidate change point $t=1,\ldots,N$, taking us from calendar time 17 May 2017 to 10 Oct 2017 in our example. With each evaluation, we are asking the question “What is the probability that this set of forecasts is actually drawn from two different distributions, one before and another after our candidate $t$?” A visualisation of a single step in the process is shown in Figure 2 where the two periods are highlighted with no shading and light gray shading, respectively. Having computed (6) for each time point, we can normalize it to derive our posterior mass function for the location of the change point. The corresponding CMF then provides weights for our aggregated forecast. For readers who might find it helpful, A provides an illustrative example of how kairosis works on simulated data with a simplified set of just three possible change points.
@@
The mass and cumulative mass functions for our running example are illustrated in Figure 3. The dominant feature here is a flurry of forecasting activity in August 2017. The forecaster clock is running fast here, increasing the concentration in chronological time of potential change points. Our posterior probability for a change point accounts for this activity and, using the Dirichlet-categorical term in (6), considers whether the distribution of forecasts actually changes here. Indeed, the distribution of forecasts appears to shift upwards and the probability of a change point is deemed to be high. The resulting CMF features a large, sudden rise with the effect that our aggregation weights also rise. (The timing of this probable change point coincides with the announcement of the handing-over, from the Trump campaign team to the US Senate Judiciary Committee, of documents relating to suspected Russian collusion in the 2016 presidential election. Although the link between the announcement and the probable change point obviously cannot be ascertained here, it does provide a plausible explanation for the data.)
@@
Despite their primary use as weighting functions, we find it useful to interpret CMFs such as those illustrated in Figure 3 as functions nonlinearly transforming chronological time to a scaled version of our notional kairos time. The concept guides our intuition when reconciling contextual information, observed forecaster data and aggregation weights.
@@
Figure 3: Kairosis weights are computed from the posterior distribution for the most recent change point. Values are calculated for the mass function and cumulative mass function for the posterior distribution over potential times for the most recent change point. They are interpolated and superposed on Subfigures 3(a) and 3(b), respectively. The cumulative mass function provides the kairosis weights for aggregating the forecasts.
@@ 2.2 Parameter Selection
Remembering that our forecaster clock is ticking with the arrival of each forecast, the change point occurrence rate $p$ can also be thought of in terms of its reciprocal $1/p$, which describes the expected number of forecasts per change point. Seen this way round, the parameter can be understood as describing the responsiveness of the population of forecasters, which in the Metaculus data remains fairly constant from question to question even though the question topics may vary considerably. In the numerical comparisons of Section 3 we have specified $101/p=$ 10, meaning that we expect a change point to motivate approximately 10 forecasters to contribute updated forecasts. Sensitivity analyses presented in Appendix B show that this choice generally leads to near-optimal Brier scores for the kairosis-aggregated forecasts and for the aggregated forecast with geometrically decaying weights. We note also that the prior distribution for the change point location tends towards a uniform distribution as $1/p$ becomes very large so that all values of $1/p\gg N$ lead to approximately the same posterior inferences.
@@
Recall that the $\alpha_{k}$ and $\alpha_{k}^{\prime}$ are pseudo-counts quantifying our prior beliefs for the proportion of forecasts falling in each bin, to either side of a putative change point. They appear in the likelihood function for the change point (equation (6)) via the quantities $\alpha_{k}+n_{k}$ and $\alpha_{k}^{\prime}+n_{k}^{\prime}$. First, consider the situation after the most recent change point. The forecasts here are understood to come from a single distribution and we expect this distribution to have low entropy, meaning that forecasts are concentrated in a small number of bins. When this is observed to be the case for a candidate change point, then our posterior probability for its being the most recent change point is increased. Small values of $\alpha_{k}^{\prime}=\alpha^{\prime}$, equal over bins, correspond to a prior that allocates high probability to such low-entropy forecast distributions. Next, consider the situation before the most recent change point. This can be more complicated, since the forecasts here could have resulted from multiple inter-change point distributions. We cannot expect them to come from a single distribution characterized by high levels of agreement among forecasters. We therefore encode the anticipated diversity of historical forecasts, which grows as that history increases in duration, by setting $\alpha_{k}=\lambda t_{\rm\small CP}$ where $\lambda>0$ is a positive scaling factor. An effect of this choice is to limit, but not ignore, the influence of distant history on our estimate for the most recent change point. In the calculations of Section 3 we have chosen to set $\lambda=0.2$. This choice is motivated by the idea that for every forecast in the distant past (i.e. before the most recent change point) we add another pseudo-forecast that is effectively spread out over the five bins. Our sensitivity analyses indicate a high level of robustness to choices of $\lambda>0.1$.

### v4

@@ Abstract
We present a new method, “kairosis”, for aggregating probability forecasts made over a time period of a single outcome determined at the end of that period. Informed by work on Bayesian change-point detection, we begin by constructing for each time during the period a posterior probability that the forecasts before and after this time are distributed differently. The resulting posterior probability mass function is integrated to give a cumulative mass function, which is used to create a weighted median forecast. The effect is to construct an aggregate in which the most heavily weighted forecasts are those which have been made since the probable most recent change in the forecasts’ distribution. Kairosis outperforms standard methods, and is especially suitable for geopolitical forecasting tournaments because it is observed to be robust across disparate questions and forecaster distributions.
@@ 1 Introduction
Geopolitical forecasting tournaments have become increasingly popular over the last decade, notable providers include the Good Judgment Project and Metaculus. A typical question from Metaculus is that of Figure 1, “Will Donald Trump be president of the USA in 2019?”. From when the question opened (May 17, 2017), forecasters submitted probability forecasts (on a scale of 00 to $1$), until the question was resolved on Feb 1, 2019, although here we show only the first seven months’ forecasts. A forecaster may make as many or as few forecasts as they wish, at any time they wish; thus we observe only the overall evolving distribution of forecasts, rather than multiple time-series of individual forecasters’ evolving beliefs.
After resolution, the forecasts are scored. If forecasts are considered “static”, taking no account of when the forecast is submitted, a simple proper probability score, such as the Brier (quadratic, Brier 1950) or Log (logarithmic, Good 1952) score, can be used. Proper scores are optimized by, and therefore incentivize forecasters to submit their best estimates of, the true probability – although propriety fails if rewards are not proportional to the score, for example if the prize goes to the overall winner (see for example Roulston 2007, Broecker and Smith 2007). But prescience is clearly valuable (Schuler et al., 2025), and Metaculus, for example, weights the score by how long it was submitted before resolution. Indeed it is clear just from a visual inspection that the distribution of forecasts is not stationary; it changes both smoothly and sharply at certain points, as does the density of forecasts submitted. The reason is obvious: news occurs much in this way, and new information is continually informing the forecasting process.
@@
However, as Himmelstein et al. (2023) note, “it is possible to envision a hybrid method … of the selection and weighting approaches”, and it is our purpose in this article to provide this. Our technique is informed by, but not equivalent to, Bayesian change point detection, and can be viewed as something like an interpolation between exponential discounting and the “most recent change point” method suggested above. It achieves this by using exponential discounting as its prior (starting) position, and then constructing for each possible change point a likelihood that this splits the forecasts into two sets, before and after the supposed change point, with distinct distributions. This gives us a posterior likelihood for every possible change point, whose probability density is then integrated to give a cumulative mass function, which is used to weight past forecasts. The practical effect of our method is to create a weighting with multiple downward (as we move into the past) steps, each corresponding to significant changes in the distribution of forecasts. At one extreme, if there is no obvious time at which the distribution of forecasts changes, the weighting stays close to the original exponential discounting of past forecasts. At the other extreme a single, obvious change point between very different distributions of forecasts is effectively a horizon, behind which old forecasts add nothing.
@@
It is clear from Figure 1 that forecasts are not made at a uniform rate in chronological time; rather there are intense bursts of activity, typically due to the publication of relevant news, interspersed by quieter periods. If $R$ forecasts are made at chronological times $t_{r}$, $r=1,\ldots,R$, then the forecast order $1,\ldots,R$ effectively creates a “forecaster clock” that dilates and magnifies the moments at which forecasts are being made at a high rate, and it is in this “forecaster time” that we search for change points. However, intense forecasting activity need not imply that the information landscape and distribution of forecasters’ beliefs are changing. To conceptualize the moments of change, we note that the ancient Greeks distinguished $\chi\rho\acute{o}\nu o\zeta$ (chronos), chronological or calendar time, from $\kappa\alpha\iota\rho\grave{o}\zeta$ (kairos), the “time” of critical moments of lived experience (Smith, 1969), and our weighting function can then be thought of as the transformation from chronos to kairos (an example of which is presented in Figure 3b). “Kairos” has a nice interpretation as “a moment of time when a prophecy was pronounced” (Tzamalikos, 2007), which fits comfortably with our conjecture that change points in forecast distributions are being driven by the arrival of new information. “Kairosis” then makes for a correspondingly concise name for our method.
@@
The article is structured as follows. In Section 2 we describe our method in detail, illustrating in Figures 13 its application to the question of the Trump presidency in 2019. In Section 3, after establishing a set of metrics and comparators, we report empirical results across 650 forecasting problems posed by Metaculus. The main part of the paper finishes with a discussion of possible developments and applications. In A we provide an additional illustrative example of computing a forecast-weighting function with our kairosis methodology in order to cement the ideas presented in the body of the paper. In B we present the preliminary results of the extension of kairosis to point forecasts of continuous variables. In C we report the results of a sensitivity analysis that accompanies the findings of Section 3.
@@ 2.1 Deriving a distribution over change point locations
Now, returning to our central example, suppose it is October 10, 2017. Our data are $R$ forecasts made at times $t_{r}$, $r=1,\ldots,R$, up to this date and we are tasked with submitting an optimal aggregated probability forecast to answer the question of the 2019 US President. It is reasonable to assume that most of the data is relevant, but to what extent is not immediately apparent. One might, for example, propose that more recent data is likely to be more informative than that from the distant past. On the assumption that critical events invalidating preceding forecasts occur independently at a constant rate over time, we are led naturally to the idea of exponentially discounting old forecasts.
We therefore begin with a prior assumption of a constant probability $p\in[0,1]$ that at least one change point-inducing event occurs between consecutive forecasts. This motivates a geometric prior distribution on the time of the last change point, $P(t_{\rm\small CP}=t_{r})=p(1-p)^{R-r}$, which follows from the idea that for $t_{r}$ to have been the last change point there must have been $R-r$ subsequent inter-forecast periods without a change point. We then update our prior distribution given knowledge of observed forecasts using Bayes’ theorem,
$P(t_{\rm\small CP}=t_{r}|\text{Forecasts})=\frac{{P(\text{Forecasts}|t_{\rm\small CP}=t_{r})}{P(t_{\rm\small CP}=t_{r})}}{P(\text{Forecasts})}.$ (3)
The next key ingredient for our methodology is thus to specify the distribution of the forecasts given a candidate change point. Inspired by work on Bayesian hypothesis testing by Holmes et al. (2015), we use the compound Dirichlet-categorical distribution to describe the number of forecasts falling in different sub-intervals (“bins”) of $[0,1]$. We use the probabilities this distribution assigns to the observed bin-counts to inform the quantity $P(\text{Forecasts}|t_{\rm\small CP}=t_{r})$ appearing in (3).
@@
Now, to propose a change point at time $t_{\rm\small CP}=t_{r}$ (that is, just after the $r$th forecast) is to suppose that forecasts up to and after $t_{r}$ follow two independent Dirichlet-categorical distributions, because an event is thought to have occurred just after $t_{r}$ that has fundamentally changed the forecasters’ (unobserved) distribution of beliefs. Of course this is never wholly true – forecasts always incorporate the previously-available information which informed forecasts before $t_{r}$ – but we now have a way of quantifying the probability that it is so. Thus, for a single change point at $t_{r}$, expression (3) becomes
$P(t_{\rm\small CP}=t_{r}|\text{Forecasts})\propto$ ${P(\text{Forecasts}|t_{\rm\small CP}=t_{r})}{P(t_{\rm\small CP}=t_{r})}$
@@
where the unprimed and primed letters denote counts $n_{k}$ and pseudo-counts $\alpha_{k}$ up to and after $t_{r}$.
@@
We then evaluate (6) for every candidate change point $r=1,\ldots,R$, taking us from calendar time 17 May 2017 to 10 Oct 2017 in our example. With each evaluation, we are asking the question “What is the probability that this set of forecasts is actually drawn from two different distributions, one up to and another after our candidate $t_{r}$?” A visualisation of a single step in the process is shown in Figure 2 where the two periods are highlighted with no shading and light gray shading, respectively. Having computed (6) for each time point, we can normalize it to derive our posterior mass function for the location of the change point. The corresponding CMF then provides weights for our aggregated forecast. For readers who might find it helpful, A provides an illustrative example of how kairosis works on simulated data with a simplified set of just three possible change points.
@@
The mass and cumulative mass functions for our running example are illustrated in Figure 3. The dominant feature here is a flurry of forecasting activity in early August 2017. The “clock” of forecaster time is running fast here, increasing the concentration in chronological time of potential change points. Our posterior probability for a change point accounts for this activity and, using the Dirichlet-categorical term in (6), considers whether the distribution of forecasts actually changes here. Indeed, it appears to shift upwards, and the probability of a change point is deemed to be high. The resulting CMF features a large, sudden rise with the effect that our aggregation weights also rise. With the vertical scale from 0 to 1 as shown, all forecasts after early August receive weights between 0.8 and 1.0, while those before are given weights less than 0.25. This relative downweighting of earlier forecasts by a factor of 3 to 4 is then kairosis’s belief of what is appropriate given its degree of confidence in the time of the most recent change point.
(It is perhaps worth noting that the timing of this probable change point coincides with the announcement of the handing-over, from the Trump campaign team to the US Senate Judiciary Committee, of documents relating to suspected Russian collusion in the 2016 presidential election. Although the link between the announcement and the probable change point obviously cannot be ascertained here, it does provide a plausible explanation for the data.)
@@
In addition to their primary use as weighting functions, we find it useful to interpret CMFs such as those illustrated in Figure 3b as functions nonlinearly transforming chronological time to a scaled version of our notional kairos time. This concept guides our intuition when reconciling contextual information, observed forecaster data and aggregation weights.
@@
Figure 3: Kairosis weights are computed from the cumulative mass function (CMF) of the posterior distribution for the most recent change point. Values are calculated for (a, upper) the mass distribution function and (b, lower) the cumulative mass function for the posterior distribution over potential times for the most recent change point. They are interpolated and superposed on Subfigures 3(a) and 3(b), respectively, with their vertical axis labels on the right of the plots. For the CMF this happens to be the same $[0,1]$ interval as for the forecasts, but the PMF is on a dissociated scale from 00 to just above its maximum value. The CMF then provides the kairosis weights for aggregating the forecasts.
@@ 2.2 Parameter Selection
Remembering that our forecaster clock is ticking with the arrival of each forecast, we can think of the prior change point occurrence rate $p$ in terms of its reciprocal $1/p$, which describes the expected number of forecasts per change point. Seen this way round, the parameter can be understood as describing the responsiveness of the population of forecasters, which in the Metaculus data remains fairly constant from question to question even though the question topics may vary considerably. In the numerical comparisons of Section 3 we have specified $101/p=$ 10, meaning that we expect a change point to motivate approximately 10 forecasters to contribute updated forecasts. Sensitivity analyses presented in C show that this choice generally leads to near-optimal Brier scores for the kairosis-aggregated forecasts and for the aggregated forecast with geometrically decaying weights – beyond the avoidance of very low values, there is no need for fine tuning of $1/p$ to achieve good results. We note also that the prior distribution for the change point location tends towards a uniform distribution as $1/p$ becomes very large so that all values of $1/p\gg R$ lead to approximately the same posterior inferences.
@@
Recall that the $\alpha_{k}$ and $\alpha_{k}^{\prime}$ are pseudo-counts quantifying our prior beliefs for the proportion of forecasts falling in each bin, to either side of a putative change point. They appear in the likelihood function for the change point (equation (6)) via the quantities $\alpha_{k}+n_{k}$ and $\alpha_{k}^{\prime}+n_{k}^{\prime}$. First, consider the situation after the most recent change point. The forecasts here are understood to come from a single distribution and we expect this distribution to have low entropy, meaning that forecasts are concentrated in a small number of bins. When this is observed to be the case for a candidate change point, then our posterior probability for its being the most recent change point is increased. Small values of $\alpha_{k}^{\prime}$, equal for all bins, correspond to a prior that allocates high probability to such low-entropy forecast distributions, and we choose $\alpha^{\prime}_{k}=1$. Next, consider the situation before the most recent change point. This can be more complicated, since the forecasts here could have resulted from multiple inter-change point distributions. We cannot expect them to come from a single distribution characterized by high levels of agreement among forecasters. We therefore encode the anticipated diversity of historical forecasts, which grows as that history increases in duration, by setting $\alpha_{k}=\lambda t_{\rm\small CP}$ where $\lambda>0$ is a positive scaling factor. An effect of this choice is to limit, but not ignore, the influence of distant history on our estimate for the most recent change point. In the calculations of Section 3 we have chosen to set $\lambda=0.2$. This choice is motivated by the idea that for every forecast in the distant past (i.e. before the most recent change point) we add another pseudo-forecast that is effectively spread out over the five bins. Our sensitivity analyses indicate a high level of robustness to choices of $\lambda>0.1$.
The overall results of the sensitivity analysis are conveyed by Figure 7. To summarize: other than the avoidance of very high values of $p$ (corresponding to a prior belief in very frequent change points) and very low values of $\lambda$ (almost no stabilization of pre-change-point forecasts) there is no need for fine tuning of $p$ or $\lambda$ in order for kairosis to achieve its results.